Maaz Munir invested 7000 part at 7% and the other at 11% if his total return for one year was 690 how much was invested at each rate
Answer by jorel1380(3719)

x+y=7000
.07x+.11y=690
7x+11y=69000
7x+7y=49000
4y=20000
y=5000
x=2000
$2000 was invested at 7%, $5000 at 11%..